OpenStudy (anonymous):

What is the y-intercept of the graph of 3x – 5y = 15 ?

OpenStudy (shane_b):

Rearrange the equation to put it in terms of y:\[y=\frac{3x-15}{5}\]Now you can simply plug that into any graphing calculator (or do it by hand) and see that the y intercept is at y=-3.

OpenStudy (lgbasallote):

you can also set x to zero then solve for y 3(0) - 5y = 15 -5y = 15 now solve for x ;)

OpenStudy (anonymous):

y=-3